I am trying to write and solve a system of equation. THe question is 3 large popcorns and 2 small drinks cost $21.00 and another customer bought 2 large popcorns and 4 small drinks for $22.00. Find the cost of a large popcorn and the cost of a small drink.
Answer by stanbon(75887)

THe question is 3 large popcorns and 2 small drinks cost $21.00 and another customer bought 2 large popcorns and 4 small drinks for $22.00. Find the cost of a large popcorn and the cost of a small drink.
Equations:
3p + 2d = 21
2p + 4d = 22
-------------------
Multiply thru the 1st equation by 2 to get:
6p + 4d = 42
2p + 4d = 22
--------------------
Subtract and solve for "p":
4p = 20
p = $5.00 (cost of one large popcorn)
------
Solve for "d":
3p + 2d = 21
3*5 + 2d = 21
15 + 2d = 21
2d = 6
d = $3.00 (cost on one drink)
==========================
Cheers,
Stan H.
=====================